Tohban Report 2011-03-09
From RHESSI Wiki
|Start Date:||12 Jan 2011|
|End Date:||19 Jan 2011|
|List all reports|
Activity is really picking up. There were three beta-gamma-delta active regions on the disk at various points during the week. Two of the three BGD regions have rotated off the disk but 11166 is at disk center. In addition, there are a number of large, bright ARs behind the east limb that will rotate on in the coming days. X-ray background is at C class level right now so the number of C class flares is likely to be more than reported below. Also, there are issues with RHESSI data from the past 24-48 hours: the SSR is not empty and these data have not yet been downloaded. Therefore the true number of RHESSI flares is not known for this week yet.
How many GOES flares occurred? Flares above B, C, M, X class were 1 51 12 0 And how many of these are listed in the RHESSI flare list? Flares above B, C, M, X class were 0 33 3 0 And how many had EXCELLENT coverage? Flares above B, C, M, X class were 0 16 0 0 There were RHESSI flares/GOES flares 348 / 57 over the time range 02-Mar-11 09-Mar-11
The SSR has a maximum of ~60% and is not emptying. Extra Weilheim and San Diago passes will be requested this week. In addition, decimation level will be increased to Active/Vigorous and the High Voltage on G9 will be decreased by 600 V (30 steps). The front segment of G9 will be turned off and the G9 rear LLD will be changed to FF. If these measure are not sufficient to lower the SSR, keeping the thin attenuator in all the time will be considered.
GAPS IN APP_ID = 154 (VC1-PMTRAS) WITH PACKET RATE LT 4 N_GAPS 1 GAP START TIME GAP END TIME GAP (SEC) 2011-03-02T00:00:00.000 -- 2011-03-02T05:05:00.000 18300.000 NO GAPS IN APP_ID = 102 (VC3-MONITOR RATES) GAPS IN APP_ID = 102 (VC3-MONITOR RATES) WITH PACKET RATE LT 30 2011-03-05T14:10:00.000 -- 2011-03-05T14:35:00.000 1500.0000 GAPS IN APP_ID = 102 (VC3-MONITOR RATES) WITH PACKET RATE LT 30 2011-03-06T10:55:00.000 -- 2011-03-06T11:30:00.000 2100.0000 2011-03-06T13:15:00.000 -- 2011-03-06T14:10:00.000 3300.0000 GAPS IN APP_ID = 1 (VC1-SOH) WITH PACKET RATE LT 30 N_GAPS 2 GAP START TIME GAP END TIME GAP (SEC) 2011-03-07T04:45:00.000 -- 2011-03-07T04:55:00.000 600.00000
Detector 9 is showing some very strange behavior which began around midnight on 05-03-11. Behavior like this has not been seen before and as yet, it remains undiagnosed. It may be associated with eclipses as the spikes are aggravated just before RHESSI passes into nighttime and the are ease after it emerges into daylight. We have tried lowering the HV by 10 steps on G9 at 1800UT on 7-march-11 to see if this helps but these data have not yet been downloaded yet. The spikes were initially triggering the attenuators to come in so detector 9 has been taken out of the attenuator formula and replaced with detector 5. Detectors 1,3,4 and 5 now determine the attenuator state. Richard Schwartz also pointed out that data drop-outs are longer in duration and more frequent when G9 is misbehaving. The gaps are 1 to 2 seconds long and can give up to 75% deadtime.
Changes to be made this week: It has been decided to lower the HV of G9 by 30 steps (600 V) before considering turning it off. Decimation will be changed to ACTIVE/VIGOROUS this evening. Front segments in detector 9 will be turned off all the time. The rear LLD in detector 9 will be changed to FF.